OpenStudy (mikee):

In Michigan, there are 2.3 lawyers for every 1000 people. The population of Detroit is 911,402. How many lawyers would you expect there to be in Detroit?

OpenStudy (anonymous):

911402/1000 = x / 2.3 x = 2096.2246

OpenStudy (mikee):

Sorry Fruitless but I am confused. What's the answer?

OpenStudy (anonymous):

The answer is x which is 2096.2246

OpenStudy (mikee):

So how many lawyers would there be?

OpenStudy (mikee):

Thank You Fruitless and Potato 1